You can jump 3 m (10 feet) off of a diving board and into a swimming pool and be uninjured. However, if you jumped from the same height onto concrete it would be very bad. Use the principles of work and energy to explain this.
Answer #1

using work-energy theroem

final velocity inboth case is given by

v^2 = 2 gh

v = sqrt ( 2 gh)

so final velocity is same in both cases

from force impulse relationship

F t =m ( v - u)

F = m ( v - u) /t

as v = constant so for change in momentum to be constant, force is inversely proportional to time

in water diver takes more time to change its momentum, hence much less value of resistive force applied by water on The diver.

but in concrete, since the change of momentum is abrupt i. e. its an instantaneous process so it takes less time, hence resistive force become much larger. as the force is a high value hence diver injured in co concrete floor.

======

Comment in case any doubt, will reply for sure.. Goodluck